# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

### v3

@@ III-E Performance Metrics
We specify the performance metrics for all estimation parameters in $\bm{\lambda}$ for the sake of evaluating Algorithm 1. For the position estimation, we take root mean square error (RMSE) as the performance index. For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
${\rm RMSE}_{u}=\sqrt{\frac{1}{M_{e}}\sum_{m=1}^{M_{e}}\|\hat{\bm{p}}_{u,m}-\bm{p}_{u}\|^{2}},$ (24)
where $M_{e}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{u,m}$ denotes the estimation of the UE position in the $m$-th trial.
@@
where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, and $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$. Using this measure, the performance of the joint CFO-PN estimation rather than each of the individual variables alone is considered.
@@ IV Hybrid Cramér-Rao Lower Bound Analysis
The HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section. This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and also serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section. The HCRLB for the estimation of $\bm{\lambda}$ is given by [34, Eq. (4.610)]
$\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,d_{Ru},\phi_{az},\phi_{el}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (27)
where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [34, Eq. (4.611)]
@@ IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}};$ (30c)
$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}};$ (30d)
$\frac{\partial\bm{\mu}}{\partial\phi_{el}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{el}},$ (30e)
@@
Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ as an example. We have
$\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$ (32)
$=P\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$
@@
Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated by
@@
Subsequently, $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ can be expressed by
@@
IV-C Derivation of the Transformed HCRLB
So far, we have derived the expression for $\mathbf{B}$, which is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$. However, when assessing the proposed estimation algorithm, we adopt the RMSE and CFO-PN joint MSE metrics introduced in Section III-E. Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$, corresponding to the two metrics. Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix by [35]
$\mathbf{HCRLB}(\bm{p}_{u},\bm{w})=\bm{\Xi}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{u}}),$ (35)
where, for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
$\bm{\Xi}(\bm{p}_{u})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{u})\right),$ (36)
with $\bm{\Xi}_{1}$ and $\bm{\Xi}_{2}(\bm{p}_{u})$ given by (37) and (38), respectively.
@@ IV-C Derivation of the Transformed HCRLB
$\bm{\Xi}_{2}(\bm{p}_{u})=\begin{bmatrix}\frac{\partial x_{u}}{\partial d_{Ru}}&\frac{\partial x_{u}}{\partial\phi_{az}}&\frac{\partial x_{u}}{\partial\phi_{el}}\\ \frac{\partial y_{u}}{\partial d_{Ru}}&\frac{\partial y_{u}}{\partial\phi_{az}}&\frac{\partial y_{u}}{\partial\phi_{el}}\\ \frac{\partial z_{u}}{\partial d_{Ru}}&\frac{\partial z_{u}}{\partial\phi_{az}}&\frac{\partial z_{u}}{\partial\phi_{el}}\\ \end{bmatrix}=\begin{bmatrix}\sin\phi_{el}\cos\phi_{az}&-d_{Ru}\sin\phi_{el}\sin\phi_{az}&d_{Ru}\cos\phi_{az}\cos\phi_{el}\\ \sin\phi_{el}\sin\phi_{az}&d_{Ru}\sin\phi_{el}\cos\phi_{az}&d_{Ru}\sin\phi_{az}\cos\phi_{el}\\ \cos\phi_{el}&0&-d_{Ru}\sin\phi_{el}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (38)
After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, the position error bound (${\rm PEB}$) for evaluating the UE position estimation and $\rm HCRLB_{CFO-PN}$ for evaluating the CFO-PN joint estimation can be given by
${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})};$ (39)
${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{0:N-1,0:N-1}).$ (40)
Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{u}$, and the RIS phase shifts, $\bm{w}$.
@@ V RIS Configuration
In this section, we further configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, i.e., PEB. According to (39), the PEB depends on the UE position, which indicates that minimizing the PEB at a specific UE position requires knowledge of that position. However, this is not possible in practice. To address this issue and enhance the algorithm’s robustness, we take the average PEB within the area of interest (AOI) as the objective function. Specifically, we randomly sample $U$ positions within the AOI and calculate their average PEB. Accordingly, the problem of interest is formulated as
$\min_{\bm{w}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\bm{w})$
$s.t.\quad|\bm{w}[r]|=1,r=1,\cdots,N_{R}.$ (41)
Problem (V) is not convex because of the non-convexity of the objective function with respect to vector $\bm{w}$, as well as the unit-module constraint.
@@
$\min_{\mathbf{W}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\mathbf{W})$
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (44)
According to (35) and (39), the optimization problem in (V) remains non-convex because of the complex objective function. To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies
$\mathbf{Z}_{u}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{u}),$ (45)
where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB. According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{u},\mathbf{W})$, (45) can be transformed to
$\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (46)
@@
$\min_{\mathbf{W},\mathbf{Z}_{1},\cdots,\mathbf{Z}_{U}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm tr}(\mathbf{Z}_{u})$
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (50)
@@
Proposition 1. $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
@@
with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.

### v4

@@ III-E Performance Metrics
We specify the performance metrics for all estimation parameters in $\bm{\lambda}$ for the sake of evaluating Algorithm 1. For the position estimation, we take root mean square error (RMSE) as the performance index. For any testing UE position $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$, ${\rm RMSE}_{\color[rgb]{0,0,0}\text{u}}$ can be expressed as
${\rm RMSE}_{\color[rgb]{0,0,0}\text{u}}=\sqrt{\frac{1}{M_{\color[rgb]{0,0,0}\text{e}}}{\color[rgb]{0,0,0}\sum_{m=0}^{M_{\text{e}}-1}}\|\hat{\bm{p}}_{{\color[rgb]{0,0,0}\text{u}},m}-\bm{p}_{\color[rgb]{0,0,0}\text{u}}\|^{2}},$ (28)
where $M_{\color[rgb]{0,0,0}\text{e}}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{{\color[rgb]{0,0,0}\text{u}},m}$ denotes the estimation of the UE position in the $m$-th trial.
@@
where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, and $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$. By focusing on joint CFO-PN estimation rather than individual variables, we ensure a more comprehensive performance evaluation. Subtracting the initial elements of $\bm{\gamma}$ and $\hat{\bm{\gamma}}$, specifically the actual PN $\bm{\theta}[0]$ and its estimate $\hat{\bm{\theta}}[0]$ at time 00, helps mitigate the influence of the initial PN $\bm{\theta}[0]$ (and its estimate $\hat{\bm{\theta}}[0]$) on the MSE analysis and resulting in a more precise evaluation of the joint CFO-PN estimation.
@@ IV Hybrid Cramér-Rao Lower Bound Analysis
HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section. This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section. HCRLB is an extension of the traditional CRLB, suitable for estimation problems with both deterministic and random unknown variables. Unlike CRLB, HCRLB exploits the prior distribution of the unknown random variables to improve analysis. Since we estimate the deterministic variables, including CFO and UE position, as well as the random variable, i.e., PN, analyzing the HCRLB allows for more accurate estimation error lower bounds for these variables.
The HCRLB for the estimate of $\bm{\lambda}$ is [43, Eq. (4.610)]
$\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,{\color[rgb]{0,0,0}d_{\text{R}}},{\color[rgb]{0,0,0}\phi_{\text{az}}},{\color[rgb]{0,0,0}\phi_{\text{el}}}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (31)
where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [43, Eq. (4.611)]
@@ IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
$\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}};$ (34c)
$\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}};$ (34d)
$\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}},$ (34e)
@@
Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ as an example. We have
$\frac{\partial\bm{\mu}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ (36)
$=P\frac{\partial\bm{h}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}},$
@@
Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated as
@@
Subsequently, $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ can be expressed as
@@
IV-C Derivation of Transformed HCRLB
So far, we have derived the expression for $\mathbf{B}$, whose inverse is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$. However, when assessing the proposed estimation algorithm, we adopt the RMSE and CFO-PN joint MSE metrics introduced in Section III-E. Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{\color[rgb]{0,0,0}\text{u}},y_{\color[rgb]{0,0,0}\text{u}},z_{\color[rgb]{0,0,0}\text{u}}\right]^{\top}$, corresponding to the two metrics. Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix [44]
$\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})=\bm{\Xi}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})\mathbf{B}^{-1}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{\color[rgb]{0,0,0}\text{u}}}),$ (39)
where, for any UE position $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$, $\mathbf{\Xi}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})$ is the transition matrix defined as
$\bm{\Xi}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})\right),$ (40)
with $\bm{\Xi}_{1}$ and $\bm{\Xi}_{2}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})$ given by (41) and (42), respectively.
@@ IV-C Derivation of Transformed HCRLB
$\bm{\Xi}_{2}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})=\begin{bmatrix}\frac{\partial x_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}&\frac{\partial x_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}&\frac{\partial x_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\\ \frac{\partial y_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}&\frac{\partial y_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}&\frac{\partial y_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\\ \frac{\partial z_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}&\frac{\partial z_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}&\frac{\partial z_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\\ \end{bmatrix}=\begin{bmatrix}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}&-{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}&{\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\\ \sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}&{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}&{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\\ \cos{\color[rgb]{0,0,0}\phi_{\text{el}}}&0&-{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (42)
After obtaining $\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})$, the position error bound (${\rm PEB}$) for the UE position estimation and $\rm HCRLB_{CFO-PN}$ for the CFO-PN joint estimation can be given by
${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})]_{N:N+2,N:N+2})};$ (43)
${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})]_{0:N-1,0:N-1}).$ (44)
Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$, and the RIS phase shifts, $\bm{w}$.
@@ V RIS Configuration
In this section, we configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, the PEB derived in Section IV. The configuration effectively enhances the performance of the algorithm proposed in Section III, and is applied before the localization process.
According to (43), the PEB depends on the UE position, which indicates that minimizing the PEB at a specific UE position requires knowledge of that position. However, this is not possible in practice. To address this issue and enhance the algorithm’s robustness, we specify the area of interest (AOI), according to the prior knowledge and practical requirements. The AOI represents the region in which we anticipate the UE of interest is situated and the high-precision localization of the UE is expected. The design objective of the RIS configuration is to achieve high expected accuracy for the localization of UEs located within the AOI, i.e., by minimizing the average PEB within the AOI. Specifically, we uniformly and randomly sample $S$ positions within the AOI, denoted as $\bm{p}_{s},s=0,1,\cdots,S-1$, and evaluate their average PEB.
This problem is transformed into a convex SDP form using the SDR method and the Schur complement. Unlike traditional approaches that rely on the classical FIM, the objective function represented by the average PEB is derived from the inverse of the BIM, $\mathbf{B}$. To establish the convexity of the transformed problem, we analytically reveal the linear relationship between $\mathbf{B}$ and the optimization variable $\mathbf{W}$.
Accordingly, the problem of interest is formulated as
$\min_{\bm{w}}\quad\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}{\rm PEB}(\bm{p}_{\color[rgb]{0,0,0}s},\bm{w})$
${\color[rgb]{0,0,0}\rm s.t.}\quad|\bm{w}[r]|=1,r={\color[rgb]{0,0,0}0,1,\cdots,}{\color[rgb]{0,0,0}N_{\text{R}}-1}.$ (45)
Problem (V) is not convex because of the non-convexity of the objective function with respect to the vector $\bm{w}$, as well as the unit-module constraint.
@@
$\min_{\mathbf{W}}\quad\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}{\rm PEB}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$
$\begin{array}[]{ll}{\color[rgb]{0,0,0}\rm s.t.}&\left[\mathbf{W}\right]_{r,r}=1,r={\color[rgb]{0,0,0}0,1},\cdots,{\color[rgb]{0,0,0}N_{\text{R}}-1},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (48)
According to (39) and (43), the optimization problem in (V) remains non-convex because of the complex objective function. To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{s}$, we define an auxiliary matrix $\mathbf{Z}_{\color[rgb]{0,0,0}s}$ that satisfies
$\mathbf{Z}_{\color[rgb]{0,0,0}s}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\mathbf{B}^{-1}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{\color[rgb]{0,0,0}s}),$ (49)
where $\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{\color[rgb]{0,0,0}s})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB. According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$, (49) can be transformed to
$\begin{bmatrix}\mathbf{Z}_{\color[rgb]{0,0,0}s}&\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{\color[rgb]{0,0,0}s})&\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (50)
@@
$\min_{\mathbf{W},\mathbf{Z}_{\color[rgb]{0,0,0}0},\cdots,\mathbf{Z}_{\color[rgb]{0,0,0}S-1}}\quad\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}{\rm tr}(\mathbf{Z}_{\color[rgb]{0,0,0}s})$
$\begin{array}[]{ll}{\color[rgb]{0,0,0}\rm s.t.}&\left[\mathbf{W}\right]_{r,r}=1,r=0,{\color[rgb]{0,0,0}1,}\cdots,{\color[rgb]{0,0,0}N_{\text{R}}-1},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}\begin{bmatrix}\mathbf{Z}_{\color[rgb]{0,0,0}s}&\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{\color[rgb]{0,0,0}s})&\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (54)
@@
Proposition 1. $\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$ is linear to $\mathbf{W}$.
@@
with $\mathbf{G}\in\mathbb{C}^{{\color[rgb]{0,0,0}N_{\text{R}}}\times N}$ and $\left[\mathbf{G}\right]_{r,n}={\color[rgb]{0,0,0}\sqrt{\rho_{{\text{aR}}}^{n,r}}}{\color[rgb]{0,0,0}\sqrt{\rho_{{\text{Ru}}}^{n,r}}}e^{-j2\pi\frac{nB}{N}({\tau_{{\color[rgb]{0,0,0}\text{a}}r}}+{\color[rgb]{0,0,0}\tau_{r\text{u}}})}$.